A 15 kg block slides on a horizontal surface with an initial speed of 12 m/s. The speed of 5.0m/s is attained after travelling a distance of 25m. (a) What is the change in kinetic energy on the box? (b) How much work was done by friction on the block?
1
Expert's answer
2021-10-05T10:06:23-0400

Given:

"m=15\:\rm kg"

"v_1=12\:\rm m/s"

"v_2=5.0\:\rm m/s"

"d=25\:\rm m"


(a) the change in kinetic energy on the box

"\Delta E_k=\frac{mv_2^2}{2}-\frac{mv_2^2}{2}\\
=\frac{15*5.0^2}{2}-\frac{15*12^2}{2}=-890\:\rm J"

(b) the work done by friction on the block

"W=\Delta E_k=-890\:\rm J"